What is the excess pressure inside a soap bubble that is 5cm in diameter, assuming 0.026Nm-1 as the surface tension of the soap solution?
1
Expert's answer
2020-12-25T14:06:10-0500

"\text{Laplace formula}"

"\Delta{P}=\sigma(\frac{1}{R_1}+\frac{1}{R_2})"

"\text{we assume that the thickness of the bubble film can be neglected:}"

"R_1 =R_2"

"\text{taking into account the fact that the soap bubble has two surfaces }"

"\text{-the inner and the outer, then:}"

"\Delta{P}=2*\sigma(\frac{1}{R}+\frac{1}{R})=\frac{4*\sigma}{R}"

"\text{where:}"

"\sigma= 0.026\text{ (surface tension)}"

"R=0.025 \text{ (radius bubble in metrs }R = \frac{d}{2})"

"\Delta{P}=\frac{4*\sigma}{R}=\frac{4*0.026}{0.025}= 4.16 \ Pa"

Answer: 4.16 Pa excess pressure inside a soap